PART 1: You paid $6,500 for the right to receive the following cash flows: years 1-4: $700, years 5-10: $900, years 11-1...

PART 1:

You paid $6,500 for the right to receive the following cash flows: years 1-4: $700, years 5-10: $900, years 11-15: $1,500. What was your rate of return?

A. 12.75%

B. 11.31%

C. 10.00%

D. 15.25%

PART 2:

Which of the four banks listed below provide you with highest effective return on your deposit. Calculate your answer to 2 decimals.

A. Bank A: 6.00% compounded annually
B. Bank B: 5.95% compounded monthly
C. Bank C: 5.90% compounded daily (365)
D. Bank D: 5.85% compounded a million times per year
Answer #1

PART 1

Home Insert Page Layout Formulas Data Re Cut Ea Copy Paste - Format Painter Clipboard Calibri 11 BI U - - Font - A A >-A- B18

your rate of return=11.31% as shown above using excel function IRR

PART 2

A. effective return=6.00%

B. effective return=(1+(5.95%/12))^12-1=6.11%

C. effective return=(1+(5.90%/365))^365-1=6.08%

D. effective return=(1+(5.85%/1000000))^1000000-1=6.02%

So highest effective return will come from  Bank B: 5.95% compounded monthly

the above is answer..